# PCGBandit: One-shot acceleration of transient PDE solvers via online-learned preconditioners

### v2

@@ 1 Introduction
Figure 1: Illustration of a learning-enhanced numerical simulation setup. In transient simulations the total wallclock cost is the sum of costs incurred at each in a sequence of discrete timesteps. Across many (but not all) practical cases, the cost at each step is dominated by solving one or more linear systems, with each solve generating data such as the solver settings used and the time needed to reach the required tolerance. By passing this feedback to a learning algorithm that configures the solver, we hope to reduce the cost of solving the linear systems by more than the cost of the additional computations performed by the learner. If so the total wallclock cost will be reduced in a one-shot manner, i.e. where learning occurs solely within the simulation itself.
We consider an alternative paradigm of learning-enhanced numerical simulation, in which the optimization algorithms have guarantees, sample complexity is zero, and correctness is inherited from classical solvers. This approach, illustrated in Figure 1, starts by identifying a repeated computation such as linear system solving that takes up a significant quantity of the cost; it then uses sequential data generated by doing these computations to speed them up on future instances, e.g. by learning better preconditioners. As shown by Khodak et al. (2024), standard adversarial bandit algorithms (Auer et al., 2002a; Zimmert and Seldin, 2021; Foster and Rakhlin, 2020) can provably learn to configure certain linear solvers using only the number of iterations as feedback; such learners are lightweight and can be directly extended to configure any solver and to use any type of feedback, including e.g. wallclock cost. Since we run no additional simulations other than the one being accelerated, and because correctness is guaranteed so long as that of the learning-free simulation is, such learning-enhanced methods can be easily compared directly to classical approaches.
@@ 2 Related work
Our contribution may be viewed as an empirical demonstration of the theoretical study by Khodak et al. (2024) of online-learned preconditioners; we implement their methodology in a widely used numerical software to show the approach’s practical utility and broad applicability beyond toy settings. Prior to that work, the underlying idea of using adversarial bandit learners to do online algorithm configuration was studied by Gagliolo and Schmidhuber (2006) and Gagliolo and Schmidhuber (2011), albeit mainly in the context of combinatorial optimization.
@@
Algorithm 1 Comparison between the classical (oblivious) adversarial bandit setting on the left and a numerical simulation code that adaptively sets solver configurations on the right, with colors denoting correspondence between specific steps. The performance of a bandit method is measured via its regret relative to the best fixed action in hindsight, whereas a numerical simulation code is evaluated according to its wallclock. In the oblivious setting these differ by an additive constant.
Input: number of actions $d$
@@
Output: regret
$\sum\limits_{t=1}^{T}\mathrm{cost}_{t}(i_{t})-\min\limits_{i\in[d]}\sum\limits_{t=1}^{T}\mathrm{cost}_{t}(i)$
Input: solver configs $1,\dots,d$
@@
Output: wallclock $\sum\limits_{t=1}^{T}\mathrm{cost}_{t}(i_{t})$
@@ 3 Methodology
We consider numerical simulations where the main computational cost is in solving a sequence $\{(\mathbf{A}_{t},\mathbf{b}_{t})\}_{t=1}^{T}$ of $T$ linear systems $\mathbf{A}_{t}\mathbf{x}=\mathbf{b}_{t}$, for SPD matrices $\mathbf{A}_{t}\in\mathbb{R}^{n\times n}$ and target vectors $\mathbf{b}_{t}\in\mathbb{R}^{n}$. Here solving refers to finding a solution vector $\mathbf{x}\in\mathbb{R}^{n}$ that satisfies a (potentially $t$-dependent) convergence criterion, e.g. $\|\mathbf{A}_{t}\mathbf{x}-\mathbf{b}_{t}\|_{2}\leq\varepsilon\|\mathbf{b}_{t}\|_{2}$ for some $\varepsilon>0$; for generality, we do not specify this formally, as OpenFOAM has a more involved criterion.222www.openfoam.com/documentation/guides/latest/doc/guide-solvers-residuals.html To find such solutions we will use conjugate gradient preconditioned with one of $d$ possible preconditioners, each of which specifies a function that constructs and applies an $\mathbf{A}_{t}$-dependent SPD transformation of the residual at each iteration of PCG. At each timestep $t\in[T]$ we select a preconditioner $i_{t}\in[d]$, solve the linear system $\mathbf{A}_{t}\mathbf{x}=\mathbf{b}_{t}$, and receive feedback $\mathrm{cost}_{t}(i_{t})\in\mathbb{R}_{\geq 0}$ corresponding to the time required for preconditioner $i_{t}$ to solve linear system $t$. The goal is to sequentially select preconditioners $\{i_{t}\}_{t=1}^{T}$ such that the total cost $\sum_{t=1}^{T}\mathrm{cost}_{t}(i_{t})$ incurred is small.
@@ 3.1 Solver configuration as an adversarial bandit problem
For simplicity, we omit two details from our setup. The first is that the sequence of linear systems is not oblivious to the choice of solver, as in practice both $\mathbf{A}_{t+1}$ and $\mathbf{b}_{t+1}$ often depend on the solution returned for the system $\mathbf{A}_{t}\mathbf{x}=\mathbf{b}_{t}$, and different solvers select different vectors in the $\varepsilon$-ball around the true solution. While this has theoretical implications in online learning, in practice the physical state of a well-specified numerical simulation should not depend strongly on the choice of solver; more concisely, we expect the realized sequence of linear systems to be $\mathcal{O}(\varepsilon)$-close to an oblivious sequence, and the tolerance $\varepsilon$ is usually small enough to safely ignore this issue.
@@ 3.2 Learning algorithm
Algorithm 2 Our PCGBandit algorithm for adaptively configuring solvers across a sequence of $T$ linear system instances. The method’s update rule differs from Tsallis-INF, as we normalize the one in Equation 3 by the average wallclock $W/t$ observed so far. The optimization algorithm used to obtain the updated distribution $\mathbf{p}_{t+1}$ is Newton’s method initialized at $\mathbf{p}_{t}$.
@@
$\mathbf{p}\leftarrow\mathbf{1}_{d}/d$
  // initialize distribution over configurations
@@
 $\mathbf{c}_{[i_{t}]}\leftarrow\mathbf{c}_{[i_{t}]}+\frac{\mathrm{cost}_{t}(i_{t})}{\mathbf{p}_{[i_{t}]}}$
@@
 $\mathbf{p}_{t+1}\leftarrow\operatorname*{arg\,min}\limits_{\mathbf{p}\in\triangle_{d}}~\langle\mathbf{p},\mathbf{c}\rangle-\frac{2W}{\sqrt{t}}\sum\limits_{i=1}^{d}\sqrt{\mathbf{p}_{[i]}}$
@@
of the empirical cost estimate (first term) regularized by the negative Tsallis entropy (second term). This minimization can be done using Newton’s method at a computational cost of $\mathcal{O}(d)$ (Zimmert and Seldin, 2021, Algorithm 2); notably this means the learning cost does not scale with the problem size $n$. We also follow Zimmert and Seldin (2021, Theorem 1) in setting $\eta_{t}=\frac{2}{\sqrt{t}}$.
When the cost of every configuration is bounded, Tsallis-INF guarantees (in expectation) that $\mathrm{Regret}_{T}=\mathcal{O}(\sqrt{dT})$, which is worst-case optimal in both the number of timesteps $T$ and the number of configurations dd. [^4] See Khodak et al. (2024) for the case where the configuration space is an interval of $\mathbb{R}$. Substituting this into Equation 2 suggests that if $T\gg d$ then our wallclock will be almost as good as if we had always used the best fixed solver. It also suggests a tradeoff between the number of configurations $d$ we can consider and the regret; in particular, more configurations will decrease the first right-hand term in Equation 2 characterizing the simulation cost, because we are minimizing over more options, but it will increase the bound of $\mathcal{O}(\sqrt{dT})$ on the regret in the second term. We explore this tradeoff empirically in Section 4.4. Note that in practice, we do not have a bound on the cost of every configuration, [^5] This issue was formally studied for by Gagliolo and Schmidhuber (2011). but we find that it suffices to normalize the first term in the update objective (3) by the average cost $\frac{1}{t}\sum_{s=1}^{t}\mathrm{cost}_{s}(i_{s})$ seen so far. See Algorithm 2 for a full specification.
Tsallis-INF is just one of many bandit algorithms that could be applied to this problem. While our choice of method was motivated by worst-case optimal guarantees, such adversarial regret bounds are often viewed as pessimistic in practical non-stochastic settings that are not (colloquially) adversarial. Future work may consider other cost estimators or other bandit algorithms, such as “best-of-both-worlds” approaches that can adapt to stationarity but retain worst-case adversarial guarantees (Xu and Zeevi, 2024).
@@ 3.3 Preconditioner configurations
We now specify the set of configurations that we learn across. The main SPD preconditioners used in OpenFOAM are diagonal-based incomplete Cholesky (DIC) and geometric agglomerated algebraic multigrid (GAMG). DIC is a no-fill incomplete Cholesky factorization computed using a recurrence for its diagonal elements (Saad, 2003) and has no configurable parameters. GAMG, on the other hand, has numerous influential hyperparameters, of which we focus on which smoother to use (either Gauss-Seidel, DIC, DIC+Gauss-Seidel, or symmetric Gauss-Seidel), the number of cells in the coarsest grid level (either 10, 100, or 1000), and whether or not to merge grid levels. The first of these is tuned because it is required, the second because it was found to be influential by Geise (2023), and the third because of its documented speedup potential on simple grids; [^6] https://www.openfoam.com/documentation/user-guide/6-solving/6.3-solution-and-algorithm-control however, other GAMG settings could also be tuned. We turn off agglomeration caching because it is not currently implemented to correctly handle varying GAMG configurations, although in-principle it could do so.
In order to enhance the search space in the case of underperformance of GAMG, we also implement a new (to OpenFOAM) preconditioner, ICTC, which is a thresholded version of incomplete Cholesky with a tuneable drop-tolerance (Brown et al., 1998). In typical incomplete Cholesky, i.e. DIC or IC(0), we only fill in the elements of the Cholesky factors of $\mathbf{A}_{t}$ corresponding to the latter’s nonzeros; by contrast, ICTC fills in more of the elements according to a threshold parameter indicating which ones to drop. A lower threshold corresponds to more of the elements being filled in, thus typically reducing the number of iterations due to a better approximation of the Cholesky factors (and thus a better approximation of $\mathbf{A}_{t}^{-1}$ by the preconditioner). This comes at a cost of a higher per-iteration cost due to the larger number of nonzero elements. While the threshold is a continuous parameter, we discretize it to choose between eight settings in the set $\{10^{-4},10^{-3.5},\dots,10^{-0.5}\}$; note that Khodak et al. (2024) show for a different preconditioner (successive over-relaxation) that a fine-enough discretization can be enough to obtain optimality over the continuous domain. Together with DIC and the GAMG settings from before this yields a search space of $d=33$ configurations.
@@ 3.4 Practical considerations
While the above largely specifies the PCGBandit algorithm, we close with a discussion of two additional practical considerations. Firstly, since most software including OpenFOAM specifies a maximum number of PCG iterations, it is possible that certain configurations do not reach the required tolerance in time and thus return an insufficiently accurate solution, thus degrading simulation correctness. To handle this we implement a backstop routine, where we rerun PCG with a default preconditioner (DIC) once the maximum number is reached; this additional cost is counted towards the cost of the original configuration.
@@
Table 1: Information about the evaluation simulations. The “wallclock” column reports the percentage of total time spent on the equation when using DIC-preconditioned conjugate gradient. Only equations accounting for at least 1% of the total wallclock are included.
Figure 2: Plots of the relative variation $\|\mathbf{A}_{t}-\mathbf{A}_{1}\|_{F}/\|\mathbf{A}_{1}\|_{F}$ of the linear system matrices $\mathbf{A}_{t}$ from the initial matrix $\mathbf{A}_{1}$ across simulation timesteps $t\in[T]$. Note that this does not fully capture the change in difficulty of solving the systems, as Figures 6 and 8 demonstrate that the per-iteration cost (i.e. the slope of the cumulative wallclock) of the Shercliff flow and fringing B-field momentum equations does change over time despite the matrix not changing significantly in Frobenius norm. By comparing with Figure 5, we see that the magnitude of the improvement due to PCGBandit does not seem to be significantly related to how much the matrices vary over time.

### v3

# One-shot acceleration of transient PDE solvers via online-learned preconditioners

@@ 1 Introduction
Figure 1: Illustration of a learning-enhanced numerical simulation setup. In transient simulations the wallclock cost is the sum of costs incurred at each in a sequence of discrete timesteps. For many practical cases, the cost at each step is dominated by solving one or more linear systems, with each solve generating data such as the solver settings used and the time needed to converge. By passing this feedback to a learning algorithm that configures the solver, we hope to reduce the cost of solving the linear systems by more than the cost of the extra computations performed by the learner. If so the total wallclock cost will be reduced in a one-shot manner, i.e. where learning occurs solely within the simulation itself.
We consider an alternative paradigm of learning-enhanced numerical simulation, in which the optimization algorithms have guarantees, the number of samples required is zero, and correctness is inherited from classical solvers. This approach, illustrated in Figure 1, starts by identifying a repeated computation such as linear system solving that takes up a significant quantity of the cost; it then uses sequential data generated by doing these computations to speed them up on future instances, e.g. by learning better preconditioners or components of preconditioners. [^1] This is useful even when a class of solvers is known to be asymptotically optimal for the setting at hand (e.g. multigrid for elliptic PDEs (Hackbusch, 1985)), as the non-asymptotic cost may still depend heavily and unpredictably on the exact configuration within that class (e.g. on the choice of smoother or interpolation operator in multigrid). For some parameters the optimal value may even depend on the computer system being used. As shown by Khodak et al. (2024), standard adversarial bandit algorithms (Auer et al., 2002b; Zimmert and Seldin, 2021; Foster and Rakhlin, 2020) can provably learn to configure certain linear solvers using only the number of iterations as feedback; such learners are lightweight and can be directly extended to configure any solver and to use any type of feedback, including e.g. wallclock cost. Since we run no additional simulations other than the one being accelerated, and because correctness is guaranteed so long as that of the learning-free simulation is, such learning-enhanced methods can be easily compared directly to classical approaches.
@@ 2 Related work
Our contribution builds upon the theoretical study by Khodak et al. (2024) of online-learned preconditioners. On a technical level, we consider multiple different learning procedures from the bandit literature and settle on a different, variance-reduced estimation technique (Zimmert and Seldin, 2021) to approximate solver costs that leads to improved empirical performance. Furthermore, we implement our approach in a widely used numerical software, demonstrating the paradigm’s practical utility and broad applicability beyond toy settings. Prior to that work, the underlying idea of using adversarial bandit learners to do online algorithm configuration was studied by Gagliolo and Schmidhuber (2006) and Gagliolo and Schmidhuber (2011), albeit mainly in the context of combinatorial optimization.
@@
Algorithm 1 Comparison between the proposed adversarial bandit approach (left) and a numerical simulation code that adaptively sets solver configurations (right), with colors denoting correspondence between specific steps. The performance of a bandit method is measured via its regret relative to the best fixed action in hindsight, whereas a simulation code is evaluated by its wallclock. These metrics differ by an additive constant.
$d\leftarrow$ number of actions
-9pt
@@
-8pt
Performance measure:
regret
$\sum\limits_{t=1}^{T}\mathrm{cost}_{t}(i_{t})-\min\limits_{i\in[d]}\sum\limits_{t=1}^{T}\mathrm{cost}_{t}(i)$ $d\leftarrow$ number of solver configs
-9pt
@@
-7pt
Performance measure:
4pt
wallclock $\sum\limits_{t=1}^{T}\mathrm{cost}_{t}(i_{t})$
@@ 3 Methodology
We consider numerical simulations where the main computational cost is in solving a sequence $\{(\mathbf{A}_{t},\mathbf{b}_{t})\}_{t=1}^{T}$ of $T$ linear systems $\mathbf{A}_{t}\mathbf{x}=\mathbf{b}_{t}$, for SPD matrices $\mathbf{A}_{t}\in\mathbb{R}^{n\times n}$ and target vectors $\mathbf{b}_{t}\in\mathbb{R}^{n}$. Here solving refers to finding a solution vector $\mathbf{x}\in\mathbb{R}^{n}$ that satisfies a (potentially $t$-dependent) convergence criterion, e.g. $\|\mathbf{b}_{t}-\mathbf{A}_{t}\mathbf{x}\|_{2}\leq\varepsilon\|\mathbf{b}_{t}\|_{2}$ for some $\varepsilon>0$; for generality, we do not specify this formally, to account for more involved criteria.333www.openfoam.com/documentation/guides/latest/doc/guide-solvers-residuals.html To find such solutions we will use conjugate gradient preconditioned with one of $d$ possible preconditioners, each of which specifies a function that constructs and applies an $\mathbf{A}_{t}$-dependent SPD transformation of the residual at each iteration of PCG. At each timestep $t\in[T]$ we select a preconditioner $i_{t}\in[d]$, solve the linear system $\mathbf{A}_{t}\mathbf{x}=\mathbf{b}_{t}$, and receive feedback $\mathrm{cost}_{t}(i_{t})\in\mathbb{R}_{\geq 0}$ corresponding to the time required for preconditioner $i_{t}$ to solve linear system $t$. The goal is to sequentially select preconditioners $\{i_{t}\}_{t=1}^{T}$ such that the total cost $\sum_{t=1}^{T}\mathrm{cost}_{t}(i_{t})$ incurred is small.
@@ 3.1 Solver configuration as an adversarial bandit problem
For simplicity, we omit two details from our setup. The first is that the sequence of linear systems is not oblivious to the choice of solver, as in practice both $\mathbf{A}_{t+1}$ and $\mathbf{b}_{t+1}$ often depend on the solution returned for the system $\mathbf{A}_{t}\mathbf{x}=\mathbf{b}_{t}$, and different solvers return different vectors in the $\varepsilon$-ball around the true solution. While this has theoretical implications in online learning, in practice the physical state of a well-specified numerical simulation should not depend strongly on the choice of solver; more concisely, we expect the realized sequence of linear systems to be $\mathcal{O}(\varepsilon)$-close to an oblivious sequence, and the tolerance $\varepsilon$ is usually small enough to safely ignore this issue.
@@ 3.2 Learning algorithm
Algorithm 2 Our PCGBandit algorithm for adaptively configuring solvers across a sequence of $T$ linear system instances. The method’s update rule differs from Tsallis-INF, as we normalize the cost estimates by the average wallclock $W/t$ observed so far. The optimization algorithm used to obtain the updated distribution $\mathbf{p}_{t+1}$ is Newton’s method initialized at $\mathbf{p}_{t}$.
@@
$\mathbf{s}\leftarrow\mathbf{0}_{d}$
// initialize variance-reduction shifts
@@
$\mathbf{p}_{1}\leftarrow\mathbf{1}_{d}/d$
// set initial distribution over configurations
@@
 $\mathbf{c}_{[i_{t}]}\leftarrow\mathbf{c}_{[i_{t}]}+\frac{\mathrm{cost}_{t}(i_{t})}{\mathbf{p}_{t[i_{t}]}}$
@@
 for index $i=1,\dots,d$ do
    if $\mathbf{p}_{t[i]}\geq 16/t$ then
       $\mathbf{s}_{[i]}\leftarrow\mathbf{s}_{[i]}+\frac{1}{2}$
       // update variance-reduction shifts
       if $i=i_{t}$ then
          $\mathbf{s}_{[i]}\leftarrow\mathbf{s}_{[i]}-\frac{1}{2\mathbf{p}_{t[i]}}$
@@
 $\mathbf{p}_{t+1}\leftarrow\operatorname*{arg\,min}\limits_{\mathbf{p}\in\triangle_{d}}~\langle\mathbf{p},\frac{\mathbf{c}}{W/t}+\mathbf{s}\rangle-\sum\limits_{i=1}^{d}\sqrt{t\mathbf{p}_{[i]}}$
@@
of the empirical cost estimate (first term) regularized by the negative Tsallis entropy (second term). This minimization can be done using Newton’s method at a computational cost of $\mathcal{O}(d)$ (Zimmert and Seldin, 2021, Algorithm 2); notably this means the learning cost does not scale with the problem size $n$.
If $\eta_{t}=\frac{2}{\sqrt{t}}$ and the cost of every configuration is bounded, then Tsallis-INF guarantees (in expectation) that $\mathrm{Regret}_{T}=\mathcal{O}(\sqrt{dT})$ (Zimmert and Seldin, 2021, Theorem 1), which is worst-case optimal in both the number of timesteps $T$ and the number of configurations dd. [^5] See Khodak et al. (2024) for the case where the configuration space is an interval of $\mathbb{R}$. Substituting this into Equation 2 suggests that if $T\gg d$ then our wallclock will be almost as good as if we had always used the best fixed solver. It also suggests a tradeoff between the number of configurations $d$ we can consider and the regret; in particular, more configurations will decrease the first right-hand term in Equation 2 characterizing the simulation cost, because we are minimizing over more options, but it will increase the bound of $\mathcal{O}(\sqrt{dT})$ on the regret in the second term. We explore this tradeoff empirically in Section 4.4.
We make two changes to the above for our final learning method, which is specified in Algorithm 2. First, unlike Khodak et al. (2024), we use reduced-variance (RV) cost estimation (Zimmert and Seldin, 2021), implemented by setting $\eta_{t}=\frac{4}{\sqrt{t}}$ and adding a shift $\mathbf{s}\in\mathbb{R}^{d}$ to $\mathbf{c}$ in the objective (3); this is to mitigate instability due to the division by $\mathbf{p}_{t[i]}$ in the default importance-weighted (IW) cost estimates. Second, we do not have a bound on the cost of every configuration, [^6] This issue was formally studied for by Gagliolo and Schmidhuber (2011). so we normalize the IW cost estimator $\mathbf{c}$ in the update objective (3) by the average cost $\frac{1}{t}\sum_{s=1}^{t}\mathrm{cost}_{s}(i_{s})$ seen so far.
Tsallis-INF is just one of many bandit algorithms that could be applied to this problem. While our description above was motivated by worst-case optimal guarantees, such adversarial regret bounds are often viewed as pessimistic in practical non-stochastic settings that are not (colloquially) adversarial. However, as shown by Zimmert and Seldin (2021), Tsallis-INF also enjoys “best-of-both-worlds” guarantees, in that it simultaneously recovers the optimal stochastic-setting regret; thus we might hope that it does well in non-stationary but still benign settings too. Future work may consider other cost estimators, alternative “best-of-both-worlds” methods (Xu and Zeevi, 2024), or bandit algorithms that provably adapt to unbounded configuration costs (Putta and Agrawal, 2022).
@@ 3.3 Preconditioner configurations
We now specify the set of configurations that we learn across. The main SPD preconditioners used in OpenFOAM are diagonal-based incomplete Cholesky (IC(0)) and geometric agglomerated algebraic multigrid (multigrid). The former is a no-fill incomplete Cholesky factorization computed using a recurrence for its diagonal elements (Saad, 2003) and has no configurable parameters. The latter, on the other hand, has numerous influential hyperparameters, of which we focus on which smoother to use (either Gauss-Seidel, IC(0), IC(0)+Gauss-Seidel, or symmetric Gauss-Seidel), the number of cells in the coarsest grid level (either 10, 100, or 1000), and whether or not to merge grid levels. The first of these is tuned because it is required, the second because it was found to be influential by Geise (2023), and the third because of its documented speedup potential on simple grids; [^7] https://www.openfoam.com/documentation/user-guide/6-solving/6.3-solution-and-algorithm-control however, other multigrid settings could also be tuned. We turn off agglomeration caching because it is not currently implemented to correctly handle varying multigrid configurations, although in-principle it could do so.
In order to enhance the search space in the case of underperformance of multigrid, we also implement a new (to OpenFOAM) preconditioner, ICT, which is the thresholded version of incomplete Cholesky with a tuneable drop-tolerance (Brown et al., 1998). In typical incomplete Cholesky, i.e. IC(0), we only fill in the elements of the Cholesky factors of $\mathbf{A}_{t}$ corresponding to the latter’s nonzeros; by contrast, ICT fills elements according to a threshold parameter indicating which ones to drop. A lower threshold corresponds to more of the elements being filled in, thus typically reducing the number of iterations due to a better approximation of the Cholesky factors (and thus a better approximation of $\mathbf{A}_{t}^{-1}$ by the preconditioner). This comes at a cost of a higher per-iteration cost due to the larger number of nonzero elements. While the threshold is a continuous parameter, we discretize it to choose between eight settings in the set $\{10^{-4},10^{-3.5},\dots,10^{-0.5}\}$; note that Khodak et al. (2024) show for a different preconditioner (successive over-relaxation) that a fine-enough discretization can be enough to obtain optimality over the continuous domain. Together with IC(0) and the multigrid settings from before this yields a search space of $d=33$ configurations.
@@ 3.4 Practical considerations
While the above largely specifies the PCGBandit algorithm, we close with a discussion of two additional practical considerations. Firstly, since most software including OpenFOAM specifies a maximum number of PCG iterations, it is possible that certain configurations do not reach the required tolerance in time and thus return an insufficiently accurate solution, thus degrading simulation correctness. To handle this we implement a backstop routine, where we rerun PCG with a default preconditioner (IC(0)) once the maximum number is reached; this additional cost is counted towards the cost of the original configuration.
@@
Table 1: Information about the evaluation simulations. The “wallclock” column reports the percentage of total time spent on the equation when using IC(0)-preconditioned conjugate gradient. Only equations accounting for at least 1% of the total wallclock are included.
Figure 2: Plots of the relative variation $\|\mathbf{A}_{t}-\mathbf{A}_{1}\|_{F}/\|\mathbf{A}_{1}\|_{F}$ of the linear system matrices $\mathbf{A}_{t}$ from the initial matrix $\mathbf{A}_{1}$ across simulation timesteps $t\in[T]$. Note that this just one indicator of change, as $A_{t}$ can vary significantly without significant changes in $\|\mathbf{A}_{t}-\mathbf{A}_{1}\|_{F}$. Indeed it does not fully capture the change in difficulty of solving the systems, as Figure 3 demonstrates that the costs of the Shercliff flow and fringing B-field momentum equations do change over time despite their matrices not changing significantly in Frobenius norm. By comparing with Figure 5, we see that PCGBandit can improve simulation performance regardless of the magnitude of variation in $\|\mathbf{A}_{t}-\mathbf{A}_{1}\|_{F}$.
Figure 3: PCG solve costs using two static baseline preconditioners, IC(0) and multigrid, plotted across the timesteps of the six simulations we consider. These plots track the mean, minimum, and maximum costs across ten simulations, demonstrating that the trajectories are largely non-stationary, undermining the main assumption of stochastic bandit methods.